OpenStudy (anonymous):

glucose is C6H12O6 a certain ammount of glucose contains 1.26 x10^21 oxygen atoms, what is the mass of glucose?

OpenStudy (australopithecus):

use avagadros number

OpenStudy (anonymous):

@Kystal ihave worked out solution but i don't understand something please help.

OpenStudy (anonymous):

@Australopithecus i did

OpenStudy (australopithecus):

Can you post it form e so I dont have to look it up

OpenStudy (anonymous):

this is someone's working..

OpenStudy (australopithecus):

nevermind I got it 6.02*10^23

OpenStudy (anonymous):

number moles O2 = No/NA= 1.26 x10^21/ 6.02 x10^23 = 0.002093mol number moles glucose: No x 1/6... my problem is that i don't understand where 1/6 came from

OpenStudy (australopithecus):

6.02*10^23atoms in 1 mole so we just use unit conversion \[\frac{1mol}{6.02*10^{23}atoms}* 1.26 x10^{21}atoms\]

OpenStudy (australopithecus):

the atoms unit cancels out and you are left with moles

OpenStudy (anonymous):

yes but thats not what the questions if asking.. @Kystal

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

k~ i'll give it a try :)

OpenStudy (anonymous):

what the question is asking *

OpenStudy (australopithecus):

Just use the formula \[Mole = \frac{Grams}{Molecular Mass}\]

OpenStudy (australopithecus):

what was the amount of moles you got?

OpenStudy (anonymous):

@Australopithecus i got 0.002093mol of O2 but we are looking for MM of glucose

OpenStudy (anonymous):

number of moles = number of oxygen atoms/ 6.022 x 10^23

OpenStudy (australopithecus):

Use gravimeter formula

OpenStudy (anonymous):

for one mol of glucose u get 6 moles of Oxigen so n(gluose) = 1/6 n(O)

OpenStudy (anonymous):

@Kystal you got that from the formula C6H12O6?

OpenStudy (anonymous):

then once u get number of moles.. u use n=m/M

OpenStudy (anonymous):

yea~ from the formula~ C6H12O6 the 6 in the formula means there are 6 moles of C and O in the compound~

OpenStudy (anonymous):

m(glucose)= 0.063g

OpenStudy (australopithecus):

\[0.00209mol \frac{180.15g/mol}{(15.9994g/mol)6}\] this will give you the moles of glucose to find the molecular mass just add up the grams per mole you find in the periodic table C has a molecular mass of 12.01g/mol O has a molecular mass of 15.9994g/mol H has a molecular mass of 1.00794g/mol add them up (12.01)6 + 12(1.00794) + (15.9994)6 = 180.15g/mol Then sub them into the formula

OpenStudy (anonymous):

@Kystal you are awesome! thank you! i was wondering where the 1/6 came from lool

OpenStudy (anonymous):

lol~ you're welcome :D

OpenStudy (anonymous):

so ig i had SO4 then for 1 mole of SO4 i have 4 moles Oxygen? @Kystal

OpenStudy (anonymous):

krystal so smart

OpenStudy (anonymous):

yeap :D

OpenStudy (anonymous):

krystal can u be my wife? :P

OpenStudy (australopithecus):

I'm getting 0.703g as the answer

OpenStudy (australopithecus):

wait nvm 0.706g

OpenStudy (australopithecus):

using gravimetric factor, assuming we have the moles of 6 oxygen atoms then, I do not see how my assumption could be wrong

OpenStudy (australopithecus):

spell check always respells gravimetric factor blah

OpenStudy (australopithecus):

Gravimeteric formula is usually pretty solid and makes sense I'm not sure I follow your solution Kystal

OpenStudy (australopithecus):

wait nevermind I see but it still doesn't explain the loss in compound in your solution did you do a lot of rounding?

OpenStudy (anonymous):

jz listen to krystal method its right

OpenStudy (australopithecus):

SNSDYoona my method is just as valid

OpenStudy (australopithecus):

As I'm calculating the ratio of mass between 6O atoms and glucose

OpenStudy (anonymous):

but ur method is too complicated.. Krystal method is a simpler way and easier to understand + efficient as it just applies the basics of chemistry

OpenStudy (anonymous):

the m(glucose) can be found in 2 ways since we already know that n(O)=2.09x10^-3 then n(C)=n(H) =2.09x10^-3 n(H)=2n(O) =4.18x10^-3 method one (calculating separately): m(glucose)=nM = (2.09x10^-3)(12.01) + (4.18x10^-3)(1.008)+(2.09x10^-3)(16) = 0.0628g method two (calculating all at once): n(glucose)=1/6n(O) =nM =3.48x10^-4 (6x12.01+ 12x1.008 + 6x16) =0.0628g

OpenStudy (australopithecus):

@SNSDYoona Graviametric factor is anything but complicated I made a mistake in my calculation \[(0.00209mol*15.9994g/mol)\frac{180.15g/mol}{15.9994g/mol)6} = 0.06275g\]